# TwinDiffusion: Enhancing Coherence and Efficiency in Panoramic Image Generation with Diffusion Models

### v3

@@ 4.2 Comparison
For reference model $\Phi$, we employ two variants: Stable Diffusion (SD) v2.0, a widely used diffusion model, and Stable Diffusion XL (SDXL) v1.0 [24], an advanced version of the SD. They respectively operate in a image space of $\mathbb{R}^{512\times 512\times 3}$ and $\mathbb{R}^{1024\times 1024\times 3}$. We align the size of image crops with the default resolution of $\Phi$, creating panoramas in ${512\times 2048}$ and ${1024\times 4096}$ correspondingly. To ensure the reliability of our results, we test 20 different prompts involving various contents and art styles, and generate 200 panoramas per prompt with 5 sets of random seeds. The results presented in Sec 4.2 are specifically obtained with a reference model of $\Phi=SD$, a crop fusion timestep of $\tau=T/2$, an adjacent control factor of $\lambda=1$, a view stride of $s_{v}=16$ and a cross stride of $s_{c}=8$. More details and comparisons about these contributing factors are thoroughly discussed in Sec 4.3 and Appendix C.
@@
Figure 4: Qualitative comparisons between MultiDiffusion (MD) and ours. We can see that our approach significantly reduces the odd joints and visible seams that commonly occur in MultiDiffusion, resulting in higher-quality panoramic images.
@@
We utilize a range of quantitative metrics focusing on the following four aspects: coherence at the intersection of crops, fidelity and diversity of the generated panoramas, compatibility with the input prompts, as well as efficiency of the optimization process. Recognizing that resizing the entire panoramic image to meet the small dimensions required by the metrics (e.g. $299^{2}$ for FID, $224^{2}$ for CLIP) could lead to significant loss of essential features and distortions, we choose to test with images cropped from panoramas at a $512^{2}$ resolution instead.
@@
Fig. 5 reflects the mean and standard deviation scores of the seven metrics for our approach and the baselines. Our TwinDiffusion stands out as the optimal method across all evaluation criteria. Coherence, the most important aspect of panoramic images, is greatly improved by our method, as reported in the first two subfigures. Additional progress can also be observed in the CLIP-aesthetic metric, implying that the increase in coherence could facilitate the compatibility and aesthetic appeal of generated results. Meanwhile, we get comparable scores in FID and IS. This keeps in line with our method’s primary focus on the seam issues, which may not have much impact on fidelity and diversity. In terms of efficiency, our method achieves better image quality without any compromise in time cost.
Figure 5: Quantitative results illustrate the mean and standard deviation scores of seven metrics comparing our approach with the baselines. Each row represents coherence, fidelity & diversity, compatibility and efficiency, respectively. Our method outperforms the others in all evaluation metrics without compromising efficiency, particularly in coherence (measured by LPIPS and DISTS), the best visual representation of panoramas.
@@ 4.3 Ablation
TwinDiffusion incorporates several key factors that contribute to its performance, including the timestep $\tau$ for introducing the Crop Fusion stage, the adjacent control factor $\lambda$ in optimization function, the view stride $s_{v}$ for cropping image patches, and the cross stride $s_{c}$ for cross sampling method.
@@
In the initial attempts at twin-image generation, we apply the Crop Fusion method throughout the entire denoising process. However, the outcomes are unsatisfactory as the generated images exhibit a distinct left-right fragmentation, where an appropriate optimization time window is needed to achieve the desirable results. We know that an earlier guidance plays a greater role in the diffusion trajectory. Thus, we decide to confine the optimization period to the early stages of the sampling process. That is, the Crop Fusion is carried out from $t:T\rightarrow{\tau}$ and then stopped during $t:\tau\rightarrow 0$, where $T$ represents the total timestep according to the specific scheduler. The relationship between $\tau$ and the stitching effect of twin images is depicted in Fig. 6. As $\tau$ decreases, we can see a gradual transition from under-optimization to over-optimization, with the best results attained by $\tau=T/2$.
Figure 6: The analysis of the proper timestep $\tau$ for introducing the Crop Fusion stage. There’s a progressive effect on $I_{2}^{*}$ with the decrease of optimization timestep $\tau$. (i) When $\tau>T/2$, it is under-optimized, as $[I_{2}^{*}]_{l}$ differs significantly from $[I_{1}]_{r}$. (ii) When $\tau=T/2$, it is well-optimized, generating the most natural and seamless $I_{2}^{*}$. (iii) When $\tau<T/2$, it is over-optimized, $I_{2}^{*}$ stays too close with $[I_{1}]_{r}$ while failing to fuse with $[I_{2}^{*}]_{r}$. The bottom lines offer two additional examples that further support our findings.
@@
The control factor $\lambda$ in Eq. 6 plays a crucial role in determining the alignment behavior of image crops. It allows us to adjust the balance between aligning closely with adjacent blocks and maintaining self-alignment, thus significantly influencing the overall quality of panoramic images. Since coherence is essential in panoramic images, we evaluate the quality using metrics such as LPIPS and DISTS, lower is better for both of them. In Fig. 7, we compare the results across a wide range of $\lambda$ values, ranging from 0.1 to 100. The findings reveal that an optimal value of $\lambda$ around 1 achieves the best balance between the two competing terms mentioned above, leading to the desired visual consistency in panoramic images.
Figure 7: Further explorations about adjacent control factor $\lambda$. It presents a comparative analysis across a wide range of $\lambda$ values, ranging from 0.1 to 100, to examine their impact on the alignment behavior and visual coherence in panoramic images based on the LPIPS-DISTS weighted metric. The results demonstrate that our approach is robust, and with the optimal $\lambda$ value around 1 achieving the best balance between reconciling with adjacent crops and maintaining self-alignment. This is reasonable and in keeping well with our objective.
@@
As discussed in Sce. 1 and Sec. 3.3, the stride of neighboring views controls the trade-off between quality and efficiency. Smaller $s_{v}$ results in better image quality but also takes a lot of time; larger $s_{v}$ can accelerate the generation speed of images but the overall quality falls. To test the effectiveness of our approach in improving this problem, we measure the aforementioned LPIPS-DISTS weighted score of generated images and the generation time over different $s_{v}$ values, making a comparison between MultiDiffusion and our TwinDiffusion with and without the two optimization method in Sec. 3. As seen in Fig.8, our method surpasses the baseline in a better quality-efficiency balance.
In particular, it demonstrates its superiority from these three aspects. (i) For quality: under the same view stride, our method, with single Crop Fusion or Cross Sampling, both reach a lower LPIPS-DISTS weighted score than the MultiDiffusion, and full TwinDiffusion reaches even lower scores, with the effects of both methods add up. (ii) For efficiency: analyzing the blue and green lines, we can observe that our Cross Sampling approach successfully enables generating comparable results within a fraction of the time required by the original MultiDiffusion, specifically achieving a reduction of N-fold when using a $N\times s_{v}$ value. (iii) For quality-efficiency trade-off: our method takes greater advantage of the efficiency gained from larger $s_{v}$ while maintaining high image quality, striking an optimal balance between quality and speed for panoramic image generation.
Figure 8: Ablation study about our method in improving the quality-speed trade-off. We make comparisons between MultiDiffusion and our TwinDiffusion (single Crop Fusion, single Cross Sampling, Crop Fusion + Cross Sampling), evaluating time and LPIPS-DISTS weighted score under $s_{v}=4,8,16,24,32,40,48$ from left to right. As seen, the effect of single Crop Fusion and Cross Sampling method is comparable, and they both outperform the MultiDiffusion baseline. With their additive effects, TwinDiffusion exhibits a more robust curve rate, demonstrating a better speed and quality trade-off.
Table 1: Here are the findings from our further exploration of the Cross Sampling strategy. Different $s_{c}$ values were tested on all sides, with the best and the second-best results marked in bold and underlined respectively. The better effects are more likely to be achieved when $s_{c}=s_{v}/2,\,s_{v}/3$, indicating that excessively fine cross strides are unnecessary. In the extreme case of $s_{c}=s_{v}/s_{v}=1$, the effect regresses to no Cross Sampling or even worse.
@@
$s_{c}=s_{v}/2$ 0.60 0.43 13.16 14.49 32.06 6.76 43.99
$s_{c}=s_{v}/3$ 0.63 0.42 14.29 14.15 31.38 6.46 45.01
$s_{c}=s_{v}/5$ 0.71 0.59 15.11 13.98 30.93 6.42 47.52
$s_{c}=s_{v}/7$ 0.76 0.62 16.00 13.98 30.25 6.48 53.60
$s_{c}=s_{v}/s_{v}$ 0.84 0.78 16.21 12.58 30.26 6.05 59.67
@@
We further investigate the influence of cross stride in the Cross Sampling method on final results. Tab. 1 provides a series of comparison results in detail. The scores obtained show a stable and consistent pattern, where the most desirable outcomes are generally achieved when $s_{c}$ is set to $s_{v}/2$ or $s_{v}/3$. This observation suggests that a finer interleaving level of sampling does not contribute to the quality scores but adds unnecessary running time.
@@ 5 Conclusion
In this paper, we have presented TwinDiffusion, an optimized framework for panoramic image generation using state-of-the-art diffusion models. Our work breaks through the existing limitations in quality and efficiency by introducing two key innovations: (i) a lightweight fusion stage to enhance coherence and (ii) an interleaved sampling method to improve generating speed. By extending this framework to wider domains, especially video synthesis and virtual reality, we can unlock new possibilities for creating dynamic and immersive visual content.
@@
We would like to express sincere gratitude to our supervisor, Yongchuan Tang, for his guidance throughout the process of this study. We also appreciate Xiaoyu Zhang, Mingyue Hu, and the entire research group for their inspiration and support. Additionally, we thank Yunhao Chen for assisting in diagnosing issues and providing crucial feedback and suggestions.
@@ Appendix C Analysis on Different Reference Models
In addition to Stable Diffusion (SD), we also conducted a series of experiments on Stable Diffusion XL (SDXL) as our reference model $\Phi$, aiming to further explore the generalization capability of our approach. Specifically, the performance of $\Phi$ was measured by internal comparisons within the reference image set. Considering that images generated by SDXL have larger dimensions of $1024^{2}$, we also crop them into $512^{2}$, which aligns with the other image sets.
Tab. A1 presents a comprehensive comparison among the methods using different reference models. All the results are obtained under the same condition as described in Sec. 4.2. As observed, even in the context of SDXL, our approach still contributes to refining the visual and semantic coherency of generated panoramas, albeit to a lesser extent than observed with the SD-based reference model.
This observation is understandable, considering the inherent attributes of SDXL as an advanced diffusion model of the conventional SD that doubles the resolution in both height and width. Despite the limitations imposed by the reference model, our approach still demonstrates its ability to improve the quality and coherence of the generated panoramas.
Table A1: Comparisons among different methods using various reference models. The best results are marked in bold, and the second-best ones are underlined. Despite the advanced performance stored in SDXL itself, our approach still enhances several metrics in panoramic image generation tasks, demonstrating its ability to improve panorama quality even using reference models with 4×4$\times$4 × resolution.
@@
SD 0.74 0.66 13.88 14.98 32.43 6.18 /
MDΦ=SD 0.68 0.39 13.46 13.20 31.55 5.81 43.92
OursΦ=SD 0.60 0.29 12.83 15.01 32.78 6.56 44.17
SDXL 0.61 0.29 11.85 16.01 33.54 6.50 /
MDΦ=SDXL 0.67 0.28 11.85 14.74 32.40 6.37 172.62
OursΦ=SDXL 0.60 0.28 11.85 15.63 32.72 6.58 174.46

### v4

@@ 4.2 Comparison
For the reference model $\Phi$, we employ two variants: the widely used diffusion model Stable Diffusion v2.0, and its advanced version Stable Diffusion XL v1.0 [22]. They respectively operate in an image space of $\mathbb{R}^{512\times 512\times 3}$ and $\mathbb{R}^{1024\times 1024\times 3}$. We align the size of crops with the default resolution of $\Phi$, creating panoramas in ${512\times 2048}$ and ${1024\times 4096}$ correspondingly. To ensure the reliability of our results, we test 20 different prompts involving various contents and art styles, and generate 200 panoramas per prompt with 5 sets of random seeds. The results presented in Sec. 4.2 are specifically obtained with a reference model of $\Phi=\text{SD}_{\text{2.0}}$, a crop fusion timestep of $\tau=T/2$, an adjacent control factor of $\lambda=1$, a view stride of $s_{v}=16$ and a cross stride of $s_{r}=8$. More details and comparisons about these contributing factors are thoroughly discussed in Sec. 4.3 and Appendix C.
@@
Figure 4: Qualitative comparisons between MultiDiffusion and ours. Our approach significantly reduces the odd joints and visible seams that commonly occur in MultiDiffusion, resulting in higher-quality panoramic images.
@@
We utilize a range of quantitative metrics focusing on the following four aspects: (i) coherence at the intersection of crops, (ii) fidelity and diversity of the generated panoramas, (iii) compatibility with the input prompts, as well as (ix) efficiency of the optimization process. Recognizing that resizing the entire panoramic image to meet the small dimensions required by the metrics (e.g. $299^{2}$ for FID, $224^{2}$ for CLIP) could lead to loss of essential features and distortions, we choose to test with images cropped from panoramas at a $512^{2}$ resolution instead.
@@
We make comparisons among the Blended Latent Diffusion [1], MultiDiffusion and our TwinDiffusion, calculating the mean and standard deviation scores of all metrics. Additionally, we measure the performance of $\Phi$ itself (i.e. the Stable Diffusion), which is measured by internal comparisons within the reference image set.
As reflected in Fig. 8, our approach stands out as the optimal method across all evaluation criteria. Coherence, the most important aspect of panoramic images, is greatly improved by our method, as reported in the first row. Additional progress can also be observed in the CLIP-aesthetic metric, implying that the increase in coherence could facilitate the compatibility and aesthetic appeal of generated results. Meanwhile, we get comparable scores in FID and IS. This keeps in line with our method’s primary focus on the seam issue, which may not have much impact on fidelity and diversity. In terms of efficiency, our method achieves better image quality without any compromise in time cost.
@@ 4.3 Ablation
TwinDiffusion incorporates several key factors that contribute to its performance, including the timestep $\tau$ for introducing the Crop Fusion stage, the adjacent control factor $\lambda$ in our optimization function, the view stride $s_{v}$ for cropping image patches, and the cross stride $s_{r}$ in the Cross Sampling method.
Figure 5: The analysis of the proper timestep $\tau$ for introducing the Crop Fusion stage. As $\tau$ decreases, a gradual transition from under-optimization to over-optimization can be observed, with the best results attained by $\tau=T/2$. The bottom rows of the figure offer two additional examples that further support our findings.
@@
In the initial attempts at twin-image generation, we apply the Crop Fusion method throughout the entire denoising process. However, the outcomes are unsatisfactory as the generated images exhibit a distinct left-right fragmentation, where an appropriate optimization time window is needed to achieve the desirable results. We know that an earlier guidance plays a greater role in the diffusion trajectory. Thus, we decide to confine the optimization period to the early stages of the sampling process. That is, the Crop Fusion is carried out from $t$: $T\rightarrow{\tau}$ and then stopped during $t$: $\tau\rightarrow 0$, where $T$ represents the total timestep of the corresponding diffusion scheduler.
The relationship between $\tau$ and the stitching effect of twin images is depicted in Fig. 5. We can see a progressive effect on $I_{2}^{*}$ with the decrease of optimization timestep $\tau$. (i) When $\tau>T/2$, it is under-optimized, as $[I_{2}^{*}]_{l}$ differs significantly from $[I_{1}]_{r}$. (ii) When $\tau=T/2$, it is well-optimized, generating the most natural and seamless $I_{2}^{*}$. (iii) When $\tau<T/2$, it is over-optimized, $I_{2}^{*}$ stays too close with $[I_{1}]_{r}$ while failing to fuse with $[I_{2}^{*}]_{r}$.
Figure 6: Further explorations about adjacent control factor $\lambda$. It presents a comparative analysis of different $\lambda$ values to study their impact on the alignment behavior and visual coherence of panoramas. The results demonstrate that our method achieves the best balance around $\lambda=1$ and is not sensitive to changes of $\lambda$ values.
@@
The control factor $\lambda$ in Eq. 6 plays a crucial role in determining the alignment behavior of image crops. It allows us to adjust the balance between aligning closely with adjacent blocks and maintaining self-alignment, thus significantly influencing the overall quality of panoramic images. Since coherence is an essential attribute in panoramas, we streamline our assessment to focus on LPIPS and DISTS metrics to represent image quality.
Figure 7: Ablation study about our method in improving the quality-speed trade-off. As seen, the effect of the single Crop Fusion and Cross Sampling method is comparable, and they both outperform the MultiDiffusion baseline. With their additive effects, TwinDiffusion exhibits a more robust rate of curve, demonstrating a better speed and quality balance.
In Fig. 6, we compare the results across a wide range of $\lambda$ values from 0.1 to 100. When $\lambda$ is between 0.1 and 80, the LPIPS-DISTS weighted score shows a very slow growth trend. It only starts to increase greatly when $\lambda$ exceeds 80. The lowest point is concentrated around 1, where our method achieves the optimal balance between the two competing terms mentioned above, leading to the desired visual consistency in panoramic images. This is reasonable and in keeping well with our objective.
@@
As discussed in Sce. 1 and Sec. 3.3, the stride of neighboring views controls the trade-off between quality and efficiency. Smaller $s_{v}$ results in better image quality but also takes a lot of time; larger $s_{v}$ can accelerate the generation speed of images but the overall quality falls. To test the effectiveness of our approach in improving this problem, we measure the generation time and the aforementioned LPIPS-DISTS weighted score of generated images under $s_{v}=4,8,16,24,32,40,48$, making a comparison between MultiDiffusion and our TwinDiffusion with and without the two optimization method.
Figure 8: Quantitative results comparing our approach with the baselines. Our method demonstrates its superiority in all aspects, particularly in coherence, the best visual representation of panoramas.
Table 1: Here are the findings from our further exploration of the Cross Sampling strategy. Different $s_{r}$ values were tested across all metrics. The better effects are more likely to be achieved when $s_{r}=s_{v}/2,\,s_{v}/3$, indicating that excessively fine cross strides are unnecessary.
@@
$s_{r}=s_{v}/2$ 0.60 0.43 13.16 14.49 32.06 6.76 43.99
$s_{r}=s_{v}/3$ 0.63 0.42 14.29 14.15 31.38 6.46 45.01
$s_{r}=s_{v}/5$ 0.71 0.59 15.11 13.98 30.93 6.42 47.52
$s_{r}=s_{v}/7$ 0.76 0.62 16.00 13.98 30.25 6.48 53.60
$s_{r}=s_{v}/s_{v}$ 0.84 0.78 16.21 12.58 30.26 6.05 59.67
As seen in Fig. 7, our method surpasses the baseline in a better quality-efficiency balance. (i) For quality: under the same view stride, both the Crop Fusion and Cross Sampling methods reach a lower LPIPS-DISTS weighted score than the MultiDiffusion, and full TwinDiffusion reaches even lower scores due to the combined effects of the two. (ii) For efficiency: analyzing the blue and green lines, we can observe that our Cross Sampling approach successfully enables generating comparable results within a fraction of the time required by the original MultiDiffusion, specifically achieving a reduction of N-fold when using a $N\times s_{v}$ value. (iii) For quality-efficiency trade-off: the slopes indicate that our method takes greater advantage of the efficiency gained from larger $s_{v}$ while maintaining high image quality, striking an optimal balance between quality and speed for panoramic image generation.
@@
We further investigate the influence of cross stride to ensure an appropriate interleaving frequency in the Cross Sampling method. In particular, $s_{r}=1$ means setting a different sampling mode per pixel, and $s_{r}=s_{v}$ is equivalent to not using our Cross Sampling method. All the results are obtained under the same condition described in Sec. 4.2.
Tab. 1 provides a series of comparisons on all sides, with the best and the second-best results marked in bold and underlined respectively. The scores show a stable and consistent pattern, where the most desirable outcomes are generally achieved when $s_{r}$ is set to $s_{v}/2$ or $s_{v}/3$. In the extreme case of $s_{r}=s_{v}/s_{v}=1$, the effect regresses to no Cross Sampling or even worse. This observation suggests that a finer interleaving level of sampling does not contribute to the quality scores but adds unnecessary running time.
@@ 5 Conclusion
In this paper, we have presented TwinDiffusion, an optimized framework for panoramic image generation using state-of-the-art diffusion models. Our work breaks through the existing limitations in quality and efficiency by introducing two key innovations: (i) a lightweight fusion stage to enhance coherence, and (ii) an interleaved sampling method to improve generating speed. By extending this promising framework to wider domains, especially virtual reality and graphic design, we can unlock new possibilities for creating dynamic and immersive visual content.
@@
We would like to express sincere gratitude to Xiaoyu Zhang, Mingyue Hu, and the entire research group for their inspiration and support. Furthermore, we thank Yunhao Chen for assisting in diagnosing issues and providing crucial feedback and suggestions.
@@ Appendix C Analysis on Different Reference Models
In addition to Stable Diffusion v2.0, we also conduct a series of experiments on Stable Diffusion XL v1.0 as our reference model $\Phi$, aiming to further explore the generalization capability of our approach. Considering that images generated by SDXL have larger dimensions of $1024^{2}$, we also crop them into $512^{2}$, which aligns with the other image sets.
Tab. A1 presents a comprehensive comparison among the methods using different reference models. As observed, even in the context of SDXL, our approach still contributes to refining the visual and semantic coherency of generated panoramas, albeit to a lesser extent than observed with the SD-based reference model. This observation is understandable, considering the inherent attributes of SDXL as an advanced diffusion model that doubles the default resolution in both height and width. Despite the limitations imposed by the reference model, our approach still enhances several metrics in panoramic image generation tasks, demonstrating its ability to improve panorama quality.
Table A1: Comparisons among different methods using various reference models. Despite the advanced performance stored in SDXL itself, our approach still demonstrates its ability to improve the quality and coherence of the generated panoramas.
@@
$\text{SD}_{\text{2.0}}$ 0.74 0.66 13.88 14.98 32.43 6.18 /
[…]
$\text{SDXL}_{\text{1.0}}$ 0.61 0.29 11.85 16.01 33.54 6.50 /
[…]